### v1

@@ Abstract
Developments in quantum technologies are increasingly important for multiple applications in quantum computing, quantum sensing, and quantum communication. However, decoherence remains a major challenge. Pulsed dynamical decoupling (DD) addresses this problem by using sequences of short pulses to nullify the average effect of the unwanted qubit-environment interactions Viola et al. (1999); Suter and Álvarez (2016). However, the pulses should ideally be instantaneous to avoid spurious effects Loretz et al. (2015), which is challenging, e.g., due to heating at low temperatures or with biological samples.
Continuous dynamical decoupling (CDD) is an alternative that does not require strong pulses but uses a continuous protecting field and has been successfully demonstrated in multiple applications Suter and Álvarez (2016); Neumann et al. (2013); Schirhagl et al. (2014); Degen et al. (2017); Balasubramanian et al. (2009); De Lange et al. (2010); Naydenov et al. (2011); Sriarunothai et al. (2018); Genov et al. (2020); Stark et al. (2018, 2017); Knowles et al. (2014); Mcguinness et al. (2011); Solomon (1959); Sage et al. (2013a); Kucsko et al. (2013); Balasubramanian et al. (2014); Hirose et al. (2012); Aiello et al. (2013); Baumgart et al. (2016). The continuous field opens an energy gap in the dressed state basis for first order noise protection, perpendicular to the field. However, the energy gap suffers from driving field fluctuations, introducing additional noise. Concatenated CDD overcomes this issue by adding more dressing fields to iteratively compensate for amplitude noise Cai et al. (2012). Major drawbacks of this approach are increased complexity and limited gate speed due to the small final energy gap. Other approaches include phase modulated double drive CDD, where the continuous first field is phase modulated to mimic the effect of a noiseless second field Cohen et al. (2017), combining continuous and pulsed DD Genov et al. (2019), or using the noise correlations between driving fields to suppress noise Salhov et al. (2024). Robust CDD schemes are also possible in multi-level systems Stark et al. (2018); Cohen et al. (2017); Timoney et al. (2011); Aharon et al. (2013, 2016); Barfuss et al. (2018); Timoney et al. (2011); Baumgart et al. (2016) but these typically involve additional overhead in terms of control parameters. This is often challenging due to, e.g., drift of the magnetic field or spatial inhomogeneity, limiting the coherence time especially at high magnetic fields, where small relative changes can lead to fast decoherence.
@@
In this work, we propose and implement experimentally continuous phased DD (CPDD), using a strong continuous field to suppress the effect of environmental noise to first order and apply phase changes at well defined time intervals that compensate both amplitude and environmental frequency noise errors to higher order. CPDD does not require short pulses and strong driving fields, which makes it particularly suitable for experiments that require low driving power due to temperature or setup limitations. The rotary echo protocol, where the periodic phases are 00 and $\pi$ Solomon (1959); Hirose et al. (2012); Aiello et al. (2013), is one known example of CPDD. However, its performance is not optimal due to frequency noise, e.g., at low driving fields. We apply phase changes that follow robust sequences, previously used for pulsed DD only Souza et al. (2012); Torosov et al. (2011); Genov et al. (2014, 2017), improving robustness to both frequency and amplitude noise.
We demonstrate quantum sensing of a signal from an external bath of hydrogen nuclear spins at low and high magnetic fields with a single NV center in diamond. We use CPDD with the phases of the popular XY8 sequence $\left(0,1,0,1,1,0,1,0\right)\pi/2$ Gullion et al. (1990); Maudsley (1986); Souza et al. (2012), labeling the new sequence continuous XY8 (CXY8). All DD protocols perform expectedly well at low field. However, only CPDD successfully detects a signal at high magnetic field as the technique is robust against amplitude errors, while imprecise control of the Rabi frequency proves detrimental for conventional CDD. Finally, we successfully combine our method with quantum heterodyne detection Schmitt et al. (2017), achieving $\mu$Hz uncertainty in the estimated frequency. Our work expands significantly the applicability of continuous dynamical decoupling, e.g., in NV centers, trapped ions or trapped atoms.
The system.—Magnetometry experiments measure the frequency and amplitude of a signal Degen et al. (2017); Hirose et al. (2012); Aiello et al. (2013); Stark et al. (2017, 2018). We demonstrate CPDD for sensing of an oscillating field (see Fig. 1). We consider the Hamiltonian
@@
where the detuning $\Delta\equiv\omega_{L}-\Omega$ and we neglect the fast oscillating terms at frequency $2\omega_{L}$, assuming $g\ll\omega_{L}$ and $\Delta\ll\omega_{L}$. When $\omega_{L}=\Omega$, known as Hartmann-Hahn resonance Hartmann and Hahn (1962), we observe oscillations in the population of the $x$ state in this basis with a frequency $g$. However, amplitude noise in $\Omega$ leads to unwanted detuning $\Delta$ that quickly causes decoherence for standard continuous dynamical decoupling (CDD) Suter and Álvarez (2016); Neumann et al. (2013); Schirhagl et al. (2014); Degen et al. (2017); Balasubramanian et al. (2009); De Lange et al. (2010); Naydenov et al. (2011); Knowles et al. (2014); Mcguinness et al. (2011); Sage et al. (2013a, b); Balasubramanian et al. (2014); Hirose et al. (2012); Aiello et al. (2013); Stark et al. (2017, 2018). Since amplitude noise is relative, its effects worsen at high magnetic fields, where stronger driving is needed to match the Larmor frequency. Achieving $\Omega=\omega_{L}$ is then practically impossible, leading to detuned oscillations, loss of contrast, and estimation bias.
@@
We analyze the effect of errors without the sensed field. CDD applies a continuous drive $\sim\sigma_{x}$ (see Eq. (2)), which we label CX. Its fidelity is $F_{\text{CX}}=1-O(\epsilon)$, where $\epsilon$ is the error in the transition probability of one $\pi$ pulse, e.g., $\epsilon=1-\cos^{2}{(\Delta t)}\approx 0.006$ when $\Delta=0.05\,\omega_{L}$up . The fidelity of CPDD with the CXY8 sequence is $F_{\text{CXY8}}=1-O(\epsilon^{3})$, which is much closer to $1$.
@@
The effect of the phase changes on the rectified signal is not trivial as only the component $\frac{g}{2}\cos{(\xi)}\sigma_{z}$ in Eq. (3) is typically unaffected by them Sup . Thus, CPDD has selectivity with respect to the signal initial phase, similarly to pulsed DD Degen et al. (2017), making it directly applicable to quantum heterodyne detection (Qdyne) Schmitt et al. (2017). Similarly to pulsed DD, the signal can produce sequence-dependent spurious harmonics Loretz et al. (2015), which can be comparable to the main harmonic because the signal is sensed by its effect during the driving. Such unwanted signals can be suppressed by randomized phase changes Sup , as in pulsed DD Wang et al. (2019).
@@
Experimental implementation.—We performed experiments on two different single Nitrogen-Vacancy centers (NV) in diamond with an isotopically enriched ${}^{12}\!$ (99.999%) layer Sup . The negative charge state of the NV center has $S=1$ triplet ground state Doherty et al. (2012). We used a permanent neodymium magnet to create an external magnetic field bias aligned with the NV axis to remove the degeneracy between the |±$|\pm 1\rangle$ states of the NV electron spin and to control the Larmor frequency of the hydrogen bath. We placed a copper wire on the diamond surface and used it to drive the $|0\rangle\rightarrow|-1\rangle$ transition resonantly, creating an effective two-level system. An arbitrary waveform generator is used for direct generation of the sequences and precise control of the phase changes. Illumination with a green laser (518518518\,518nm), used for initialization and readout of the NV spin state, reveals a spin dependent fluorescence and favorable relaxation into state $|0\rangle$ L’́uhmann et al. (2011); Harrison et al. (2006). All experimental details can be found in Sup .
@@
First, we demonstrate the robustness of CPDD for sensing the amplitude $g\approx 2\pi~{}11.7$ kHz of a weak oscillating field. The results from a numerical simulation and experiments are shown in Fig. 2. When the Rabi frequency of the driving field is equal to the frequency of the sensed field, i.e., $\Omega=\omega_{L}=2\pi~{}8$ MHz, we observe oscillations of the population of state $|0\rangle$ for both CX and CXY8 with $P=\cos(\Theta(t)/2)$, where $\Theta=gt$. The oscillations are quickly lost with CX even with small $\Delta\approx 2\pi~{}20$ kHz, i.e., of the order of the coupling strength $g$. On the contrary, with CXY8 the robustness is improved by more than twenty times, as we observe oscillations even for $\Delta=2\pi~{}400$ kHz Sup . The experimental results match the simulation very well with the worse performance of CDD in the experiment mainly due to amplitude fluctuations, which are not included in the simulation.
@@
Figure 3: (a) Sketch of a shallow NV center at depth $d$. It interacts with the hydrogen bath on the diamond surface within a volume of roughly $d^{3}$ Müller et al. (2014). (b) A simplified level scheme of the NV center ground state. When $\Omega=\omega_{L}$ the interaction between the NV and the hydrogen bath leads to a dip in fluorescence for CX. The resonance condition for XY8 (CXY8) is $\tau=\pi/\omega_{L}$ ($T=\pi/\omega_{L}$), see Fig. 1. Experimental results for (c) DD spectra in the low field regime. The larger dip for CX is due to its longer interaction time. The label (C)XY8-$N$ shows that the sequence is repeated $N$ times. CXY8-16-4 denotes an CXY8-16 sequence where we did not vary $\Omega$ continuously as we varied $T$ but it took only four values $2\pi\{1.9,2,2.1,2.2\}$ MHz, demonstrating CXY8 robustness. DD order scan (d) off-resonance and (e) on resonance with the Larmor frequency at low field; (f) DD spectrum in the high field regime. No contrast is observed for CX even though a longer interaction time is used than for CXY8. DD order scan (g) off-resonance and (h) on resonance with the Larmor frequency at high field. In (h) we included the off-resonance CX plot due to negligible difference from CX on-resonance, as shown in (f). The respective interaction times used in (c) and (f) are marked in plots (d,e) and (g,h) with gray (CX) and red (CXY8), vertical, dashed lines.
@@
In a second experiment, we detected a nano-NMR signal from the external hydrogen bath of the immersion oil with a shallow (depth $\approx 10$nm) NV center (Fig.3(a)). We performed the experiment at two bias fields, parallel to the NV axis: $B_{z}=(485\pm 1)$G (low field) to polarize the inherent nitrogen nuclear spin, with the Larmor frequency of the hydrogen nuclear spins $\omega_{L}\approx 2\pi~{}2.065$ MHz. Second, we applied $B_{z}=(1882\pm 3)$G (high field) where the Larmor frequency is $\omega_{L}\approx 2\pi~{}8.005$ MHz. The inherent nitrogen nuclear spin is not polarized then but the effect is negligible due to the high Rabi frequency and the decoupling sequences Vetter et al. (2022). At low field, the Rabi frequency was roughly $142\pi~{}$ 14 MHz for pulsed DD. Hence, $\frac{\omega_{L}}{\Omega}\approx 0.14$, giving a good approximation for instantaneous pulses. Figure 3(b) shows the level scheme and the Hartmann-Hahn resonance condition for CDD. Figure 3(c-e) shows the results of three different sequences in the low field regime. In all cases the hydrogen dip is clearly visible with a difference in dip size as expected from theory Sup .
@@
At high field, the necessary shorter time spacing between pulses limited the microwave Rabi frequency to around $122\pi~{}$ 12 MHz due to heating. Therefore, we could not assume instantaneous pulses for pulsed DD. Thus, we only performed continuous drive sequences at high field. Figure 3(f-h) shows a dip in fluorescence due to hydrogen, which was visible only with CXY8, even though we used a longer interaction time for standard CX. This is due to the better frequency selectivity of CXY8 as its control parameter is the time interval of the phase changes $T$. In contrast, the CX requires $\omega_{L}=\Omega$, depending on the microwave Rabi frequency, which suffers from amplitude noise, mainly due to power fluctuations Fedder et al. (2011). In our experiments, the vertical resolution of the arbitrary waveform generator is $8-bit$, resulting in a $30-40$ kHz error in the Rabi frequency, making it practically impossible to lock the driving frequency on a signal with a frequency width of about $40$ kHz. On the contrary, the CXY8 resonance condition allows a Rabi frequency error of more than ±$\pm 5\%$ around $\omega_{L}$ (see Fig. 2 and Sup ). Thus, the frequency selectivity is determined by the time digitization of the driving signal, i.e., $65$ GHz, resulting in a much smaller error than the experimental requirements. We note that the broader width of the proton dip in the high field experiment is due to the larger magnetic noise and diffusion broadening Staudenmaier et al. (2022).
@@
Discussion.—The sensitivity of CPDD and pulsed DD $\sim\alpha^{-1}$ differ by their signal attenuation factor $\alpha=1/2$ for CPDD ($\alpha=2/\pi$ for pulsed DD), see Eq. (3). The dynamic range of CPDD is similar to the one of dressed states methods e.g., continuous dynamical decoupling Degen et al. (2017), and is 0.1-100 MHz for electronic spins or superconducting qubits Degen et al. (2017) with the additional advantage of improved robustness. It can be expanded by incorporating a second driving field to sense GHz frequencies, similar to double drive continuous decoupling Cai et al. (2012); Salhov et al. (2024). The sensitivities can be improved even further by using NV centers with longer coherence times, e.g., in diamond with isotopically enriched ${}^{12}\!$ Osterkamp et al. (2015), enhancing photon collection efficiency Hadden et al. (2010); Siyushev et al. (2010); Sup or with ensembles of NV centers Barry et al. (2020). Furthermore, CPDD can be modified to perform robust quantum gates. As their speed is limited by the frequency difference of the eigenstates of the last dressed qubit, CPDD will be faster than alternative continuous robust schemes like concatenated CDD Cai et al. (2012).
@@
Acknowledgment.— This work was supported by the German Federal Ministry of Research (BMBF) by future cluster QSENS and projects DE-Brill (No. 13N16207), SPINNING, DIAQNOS (No. 13N16463), quNV2.0 (No. 13N16707), QR.X and Quamapolis (No. 13N15375), DLR via project QUASIMODO (No.50WM2170), Deutsche Forschungsgemeinschaft (DFG) via Projects No. 386028944, No. 445243414, and No. 387073854 and Excellence Cluster POLiS, BMBF via project EXTRASENS (No. 13N16935), European Union’s HORIZON Europe program via projects QuantERA II (No. 101017733), QuMicro (No. 101046911), SPINUS (No. 101135699), CQuENS (No. 101135359), QCIRCLE (No. 101059999), and FLORIN (No. 101086142), European Research Council (ERC) via Synergy grant HyperQ (No. 856432), Carl-Zeiss-Stiftung via the Center of Integrated Quantum Science and technology (IQST) and project Utrasens-Vir, as well as the Baden-Württemberg Foundation. A.R. acknowledges the support of ERC grant QRES, project number 770929, Quantera grant MfQDS, and the Schwartzmann university chair. D. L., G. T. G., and N. S. contributed equally to this work.
@@
In order to introduce the mechanism of CPDD, we first describe the theory of sensing by continuous dynamical decoupling (CDD) Suter and Álvarez (2016); Schirhagl et al. (2014); Degen et al. (2017); Balasubramanian et al. (2009); De Lange et al. (2010); Naydenov et al. (2011); Stark et al. (2018, 2017); Viola et al. (1999); Neumann et al. (2013); Knowles et al. (2014); Mcguinness et al. (2011); Sage et al. (2013a, b); Balasubramanian et al. (2014); Hirose et al. (2012); Aiello et al. (2013); Stark et al. (2018).
@@
where $\omega_{0}$ is the Bohr transition frequency, $\Omega$ and $\omega$ are the Rabi frequency and carrier (angular) frequency of the driving field with a phase $\phi$, and $\omega_{L}=\Omega+\Delta$ is the (angular) frequency of the signal (e.g., equal to the Larmor frequency of the proton spins in the hydrogen bath in our nano-NMR experiment) and $\Delta$ is its detuning from the Rabi frequency $\Omega$, $g$ is the signal amplitude, and $\xi$ is its unknown (random) phase. In the case of an artificial signal, the latter could be generated by an oscillating magnetic field $B\cos{(\omega_{L}t+\xi)}$. The value of $g=\gamma_{\text{NV}}B/2$, where $\gamma_{\text{NV}}$ is the gyromagnetic ratio of the NV center electron spin, when the detector qubit consists of the $|0\rangle$ and $|-1\rangle$ ground states of the electron spin of an NV center, as in our case (or is a spin $1/2$ system, in general). In comparison, $g=\gamma_{\text{NV}}B$ when the detector qubit consists of the |±$|\pm 1\rangle$ ground states of the NV center on the so called double-quantum transition Myers et al. (2017).
@@
In summary, the possibility to apply arbitrary phase changes in CPDD leads to selectivity with respect to the initial phase of the signal and the selectivity condition is similar as for standard pulsed DD schemes Degen et al. (2017). Finally, we note that we can choose the “good” initial phase $\xi$ of the signal by shifting the whole CPDD sequence in time. We only need to change the duration of the first pulse period to satisfy $\omega_{L}t_{1}+\xi=0~{}(\text{mod}~{}\pi)$ and apply the subsequent phase changes at intervals of $T=\pi k/\omega_{L},~{}k\in\mathbb{N}$. The phase selectivity of the protocol implies that, unlike standard CDD, it is directly applicable in synchronized sensing schemes, e.g., the Qdyne protocol, which are typically limited only by the coherence time of the signal or a classical clock Schmitt et al. (2017).
@@
The propagator that characterizes the evolution of the qubit can in general be parameterized by Genov et al. (2014, 2017); Bruns et al. (2018)
@@
where $p\equiv 1-\epsilon$ is the transition probability, i.e., the probability that the qubit will be transferred to state $|-1\rangle$ if it was initially in state $|0\rangle$, $\epsilon\in[0,1]$ is the unknown error in the transition probability, $\alpha$ and $\beta$ are unknown phases. In case of a perfect pulse, the transition probability becomes $p=1$ and $\epsilon=0$. However, this is often not the case, e.g., due to frequency or amplitude drifts or inhomogeneity, which make $\epsilon\neq 0$. For example, an amplitude error of $\Delta/\omega_{L}=(\omega_{L}-\Omega)/\omega_{L}=5\,\%$ from the target Rabi frequency of $\omega_{L}$ (Hartmann-Hahn resonance) leads to $\epsilon=1-\cos^{2}\left(\Delta t/2\right)\approx 0.006$ for a target $\omega_{L}t=\pi$ pulse, assuming $\delta=0$. Such errors can be compensated by applying phased sequences of pulses, where the phases of the subsequent pulses are chosen to cancel the errors of the individual pulses up to a certain order Genov et al. (2014, 2017); Bruns et al. (2018).
@@
and the phases $\phi_{k}$ of the individual pulses can be used as control parameters to achieve a robust performance. We can evaluate the latter by considering the fidelity Souza et al. (2012); Genov et al. (2017)

### v2

@@ Abstract
Developments in quantum technologies are increasingly important for multiple applications in quantum computing, quantum sensing, and quantum communication. However, decoherence remains a major challenge. Pulsed dynamical decoupling (DD) addresses this problem by using sequences of short pulses to nullify the average effect of the unwanted qubit-environment interactions [1, 2]. However, the pulses should ideally be instantaneous to avoid spurious effects [3], which is challenging, e.g., due to heating at low temperatures or with biological samples.
Continuous dynamical decoupling (CDD) is an alternative that does not require strong pulses but uses a continuous protecting field and has been successfully demonstrated in multiple applications [2, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The continuous field opens an energy gap in the dressed state basis for first order noise protection, perpendicular to the field. However, the energy gap suffers from driving field fluctuations, introducing additional noise. Concatenated CDD overcomes this issue by adding more dressing fields to iteratively compensate for amplitude noise [23]. Major drawbacks of this approach are increased complexity and limited gate speed due to the small final energy gap. Other approaches include phase modulated double drive CDD, where the continuous first field is phase modulated to mimic the effect of a noiseless second field [24], combining continuous and pulsed DD [25], or using the noise correlations between driving fields to suppress noise [26]. Robust CDD schemes are also possible in multi-level systems [12, 24, 27, 28, 29, 30, 27, 22] but these typically involve additional overhead in terms of control parameters. This is often challenging due to, e.g., drift of the magnetic field or spatial inhomogeneity, limiting the coherence time especially at high magnetic fields, where small relative changes can lead to fast decoherence.
@@
In this work, we propose and implement experimentally continuous phased DD (CPDD), using a strong continuous field to suppress the effect of environmental noise to first order and apply phase changes at well defined time intervals that compensate both amplitude and environmental frequency noise errors to higher order. CPDD does not require short pulses and strong driving fields, which makes it particularly suitable for experiments that require low driving power due to temperature or setup limitations. The rotary echo protocol, where the periodic phases are 00 and $\pi$ [16, 20, 21], is one known example of CPDD. However, its performance is not optimal due to frequency noise, e.g., at low driving fields. Windowless sequences of $\pi/2$ pulses and composite pulses have also been applied for decoupling and refocusing of pulse errors [31, 32, 33, 34, 35]. In this letter, we apply phase changes for continuous decoupling that follow robust sequences of $\pi$ pulses, previously used for pulsed DD only [36, 37, 35, 38], improving robustness to both frequency and amplitude noise.
We demonstrate quantum sensing of a signal from an external bath of hydrogen nuclear spins at low and high magnetic fields with a single NV center in diamond. We use CPDD with the phases of the popular XY8 sequence $\left(0,1,0,1,1,0,1,0\right)\pi/2$ [39, 40, 36], labeling the new sequence continuous XY8 (CXY8). All DD protocols perform expectedly well at low field. However, only CPDD successfully detects a signal at high magnetic field as the technique is robust against amplitude errors, while imprecise control of the Rabi frequency proves detrimental for conventional CDD. Finally, we successfully combine our method with quantum heterodyne detection [41], achieving $\mu$Hz uncertainty in the estimated frequency. Our work expands significantly the applicability of continuous dynamical decoupling, e.g., in NV centers, trapped ions or trapped atoms.
The system.—Magnetometry experiments measure the frequency and amplitude of a signal [6, 20, 21, 13, 12]. We demonstrate CPDD for sensing of an oscillating field (see Fig. 1). We consider the Hamiltonian
@@
where the detuning $\Delta\equiv\omega_{L}-\Omega$ and we neglect the fast oscillating terms at frequency $2\omega_{L}$, assuming $g\ll\omega_{L}$ and $\Delta\ll\omega_{L}$. Due to this rotating-wave approximation, the detection signal effective amplitude is $g/2\equiv\kappa g$, i.e., it is attenuated by a factor $\kappa=1/2$ in comparison to Eq. (2). We note that attenuation is slightly higher than that of pulsed DD where $\kappa_{\text{pulsed}}=2/\pi$ [6]. When $\omega_{L}=\Omega$, known as Hartmann-Hahn resonance [42], we observe oscillations in the population of the $x$ state in this basis with a frequency $g$. However, amplitude noise in $\Omega$ leads to unwanted detuning $\Delta$ that quickly causes decoherence for standard continuous dynamical decoupling (CDD) [2, 4, 5, 6, 7, 8, 9, 14, 15, 17, 43, 19, 20, 21, 13, 12]. Since amplitude noise is relative, its effects worsen at high magnetic fields, where stronger driving is needed to match the Larmor frequency. Achieving $\Omega=\omega_{L}$ is then practically impossible, leading to detuned oscillations, loss of contrast, and estimation bias.
@@
The effect of the phase changes on the rectified signal is not trivial as only the component $\frac{g}{2}\cos{(\xi)}\sigma_{z}$ in Eq. (3) is typically unaffected by them [44]. Thus, CPDD has selectivity with respect to the signal initial phase, similarly to pulsed DD [6], making it directly applicable to quantum heterodyne detection (Qdyne) [41]. Similarly to pulsed DD, the signal can produce sequence-dependent spurious harmonics [3], which can be comparable to the main harmonic because the signal is sensed by its effect during the driving. Such unwanted signals can be suppressed by randomized phase changes [44], as in pulsed DD [53].
@@
Experimental implementation.—We performed experiments on two different single Nitrogen-Vacancy centers (NV) in diamond with an isotopically enriched ${}^{12}\!$ (99.999%) layer [44]. The negative charge state of the NV center has $S=1$ triplet ground state [54]. We used a permanent neodymium magnet to create an external magnetic field bias aligned with the NV axis to remove the degeneracy between the |±$|\pm 1\rangle$ states of the NV electron spin and to control the Larmor frequency of the hydrogen bath. We placed a copper wire on the diamond surface and used it to drive the $|0\rangle\rightarrow|-1\rangle$ transition resonantly, creating an effective two-level system. An arbitrary waveform generator is used for direct generation of the sequences and precise control of the phase changes. Illumination with a green laser (518518518\,518nm), used for initialization and readout of the NV spin state, reveals a spin dependent fluorescence and favorable relaxation into state $|0\rangle$ [55, 56]. All experimental details can be found in [44].
@@
First, we demonstrate the robustness of CPDD for sensing the amplitude $g\approx 2\pi~{}11.7$ kHz of a weak oscillating field. The results from a numerical simulation and experiments are shown in Fig. 2. When the Rabi frequency of the driving field is equal to the angular frequency of the sensed field, i.e., $\Omega=\omega_{L}=2\pi~{}8$ MHz, we observe oscillations of the population of state $|0\rangle$ for both CX and CXY8 with $P=\cos(\Theta(t)/2)$, where $\Theta=gt$. The oscillations are quickly lost with CX even with small $\Delta\approx 2\pi~{}20$ kHz, i.e., of the order of the coupling strength $g$. On the contrary, with CXY8 the robustness is improved by more than twenty times, as we observe oscillations even for $\Delta=2\pi~{}400$ kHz [44]. The experimental results match the simulation very well with the worse performance of CDD in the experiment mainly due to amplitude fluctuations, which are not included in the simulation.
@@
Figure 3: (a) Sketch of a shallow NV center at depth $d$. It interacts with the hydrogen bath on the diamond surface within a volume of roughly $d^{3}$ [57]. (b) A simplified level scheme of the NV center ground state. When $\Omega=\omega_{L}$ the interaction between the NV and the hydrogen bath leads to a dip in fluorescence for CX. The resonance condition for XY8 (CXY8) is $\tau=\pi/\omega_{L}$ ($T=\pi/\omega_{L}$), see Fig. 1. Experimental results for (c) DD spectra in the low field regime. The larger dip for CX is due to its longer interaction time. The label (C)XY8-$N$ shows that the sequence is repeated $N$ times. CXY8-16-4 denotes a CXY8-16 sequence where we did not vary $\Omega$ continuously as we varied $T$ but it took only four values $2\pi\{1.9,2,2.1,2.2\}$ MHz, demonstrating CXY8 robustness. DD order scan (d) off-resonant and (e) resonant with the Larmor frequency at low field; (f) DD spectrum in the high field regime. No contrast is observed for CX even though a longer interaction time is used than for CXY8. DD order scan (g) off-resonant and (h) resonant with the Larmor frequency at high field. In (h) we included the off-resonant CX plot due to negligible difference from CX on resonance, as shown in (f). The respective interaction times used in (c) and (f) are marked in plots (d,e) and (g,h) with gray (CX) and red (CXY8), vertical, dashed lines.
@@
In a second experiment, we detected a nano-NMR signal from the external hydrogen bath of immersion oil with a shallow (depth $\approx 10$nm) NV center (Fig.3(a)). We performed the experiment at two bias fields, parallel to the NV axis: $B_{z}=(485\pm 1)$G (low field) to polarize the inherent nitrogen nuclear spin, with the Larmor frequency of the hydrogen nuclear spins $\omega_{L}\approx 2\pi~{}2.065$ MHz. Second, we applied $B_{z}=(1882\pm 3)$G (high field) where the Larmor frequency is $\omega_{L}\approx 2\pi~{}8.005$ MHz. The inherent nitrogen nuclear spin is not polarized then but the effect is negligible due to the high Rabi frequency and the decoupling sequences [58]. At low field, the Rabi frequency was roughly $142\pi~{}$ 14 MHz for pulsed DD. Hence, $\frac{\omega_{L}}{\Omega}\approx 0.14$, giving a good approximation for instantaneous pulses. Figure 3(b) shows the level scheme and the Hartmann-Hahn resonance condition for CDD. Figure 3(c-e) shows the results of three different sequences in the low field regime. In all cases the hydrogen dip is clearly visible with a difference in dip size as expected from theory [44].
@@
At high field, the necessary shorter time spacing between pulses limited the microwave Rabi frequency to around $122\pi~{}$ 12 MHz due to heating. Therefore, we could not assume instantaneous pulses for pulsed DD. Thus, we only performed continuous drive sequences at high field. Figure 3(f-h) shows a dip in fluorescence due to hydrogen, which was visible only with CXY8, even though we used a longer interaction time for standard CX. This is due to the better frequency selectivity of CXY8 as its control parameter is the time interval of the phase changes $T$. In contrast, the CX requires $\omega_{L}=\Omega$, depending on the microwave Rabi frequency, which suffers from amplitude noise, mainly due to power fluctuations [59]. In our experiments, the vertical resolution of the arbitrary waveform generator is $8-bit$, resulting in a $30-40$ kHz error in the Rabi frequency, making it practically impossible to lock the driving frequency on a signal with a frequency width of about $40$ kHz. On the contrary, the CXY8 resonance condition allows a Rabi frequency error of more than ±$\pm 5\%$ around $\omega_{L}$ (see Fig. 2 and [44]). Thus, the frequency selectivity is determined by the time digitization of the driving signal, i.e., $65$ GHz, resulting in a much smaller error than the experimental requirements. We note that the broader width of the proton dip in the high field experiment is due to the larger magnetic noise and diffusion broadening [60].
@@
Discussion.—The sensitivity of CPDD and pulsed DD differ by their signal attenuation factor $\kappa=1/2$ and $\kappa_{\text{pulsed}}=2/\pi$, respectively (see Eq. (3) and [6, 44]). The dynamic range of CPDD is similar to the one of dressed states methods e.g., continuous dynamical decoupling [6], and is 0.1-100 MHz for electronic spins or superconducting qubits [6] with the additional advantage of improved robustness. It can be expanded by incorporating a second driving field to sense GHz frequencies, similar to double drive continuous decoupling [23, 26]. The sensitivities can be improved even further by engineering of the host diamond properties to obtain NV centers with longer coherence times [62, 63], enhancing photon collection efficiency [64, 65, 44] or with ensembles of NV centers [66]. Furthermore, CPDD can be modified to perform robust quantum gates. As their speed is limited by the frequency difference of the eigenstates of the last dressed qubit, CPDD will be faster than alternative continuous robust schemes like concatenated CDD [23].
@@
Acknowledgment.— This work was supported by the German Federal Ministry of Research (BMBF) by future cluster QSENS and projects DE-Brill (No. 13N16207), SPINNING, DIAQNOS (No. 13N16463), quNV2.0 (No. 13N16707), QR.X and Quamapolis (No. 13N15375), DLR via project QUASIMODO (No.50WM2170), Deutsche Forschungsgemeinschaft (DFG) via Projects No. 386028944, No. 445243414, and No. 387073854 and Excellence Cluster POLiS, BMBF via project EXTRASENS (No. 13N16935), European Union’s HORIZON Europe program via projects QuantERA II (No. 101017733), QuMicro (No. 101046911), SPINUS (No. 101135699), CQuENS (No. 101135359), QCIRCLE (No. 101059999), and FLORIN (No. 101086142), European Research Council (ERC) via Synergy grant HyperQ (No. 856432), Carl-Zeiss-Stiftung via the Center of Integrated Quantum Science and technology (IQST) and project Utrasens-Vir, as well as the Baden-Württemberg Foundation. A.R. acknowledges the support of ERC grant QRES, project number 770929, Quantera grant MfQDS, ISF, the Schwartzmann university chair and the Israeli Innovation Authority under the project “Quantum Computing Infrastructures”.
@@
In order to introduce the mechanism of CPDD, we first describe the theory of sensing by continuous dynamical decoupling (CDD) [2, 5, 6, 7, 8, 9, 12, 13, 1, 4, 14, 15, 17, 43, 19, 20, 21, 12].
@@
where $\omega_{0}$ is the Bohr transition frequency, $\Omega$ and $\omega$ are the Rabi frequency and carrier (angular) frequency of the driving field with a phase $\phi$, and $\omega_{L}=\Omega+\Delta$ is the (angular) frequency of the signal (e.g., equal to the Larmor frequency of the proton spins in the hydrogen bath in our nano-NMR experiment) and $\Delta$ is its detuning from the Rabi frequency $\Omega$, $g$ is the signal amplitude, and $\xi$ is its unknown (random) phase. In the case of an artificial signal, the latter could be generated by an oscillating magnetic field $B\cos{(\omega_{L}t+\xi)}$. The value of $g=\gamma_{\text{NV}}B/2$, where $\gamma_{\text{NV}}$ is the gyromagnetic ratio of the NV center electron spin, when the detector qubit consists of the $|0\rangle$ and $|-1\rangle$ ground states of the electron spin of an NV center, as in our case (or is a spin $1/2$ system, in general). In comparison, $g=\gamma_{\text{NV}}B$ when the detector qubit consists of the |±$|\pm 1\rangle$ ground states of the NV center on the so called double-quantum transition [49].
@@
In summary, the possibility to apply arbitrary phase changes in CPDD leads to selectivity with respect to the initial phase of the signal and the selectivity condition is similar as for standard pulsed DD schemes [6]. Finally, we note that we can choose the “good” initial phase $\xi$ of the signal by shifting the whole CPDD sequence in time. We only need to change the duration of the first pulse period to satisfy $\omega_{L}t_{1}+\xi=0~{}(\text{mod}~{}\pi)$ and apply the subsequent phase changes at intervals of $T=\pi k/\omega_{L},~{}k\in\mathbb{N}$. The phase selectivity of the protocol implies that, unlike standard CDD, it is directly applicable in synchronized sensing schemes, e.g., the Qdyne protocol, which are typically limited only by the coherence time of the signal or a classical clock [41].
@@
The propagator that characterizes the evolution of the qubit can in general be parameterized by [35, 38, 45]
@@
where $p\equiv 1-\epsilon$ is the transition probability, i.e., the probability that the qubit will be transferred to state $|-1\rangle$ if it was initially in state $|0\rangle$, $\epsilon\in[0,1]$ is the unknown error in the transition probability, $\alpha$ and $\beta$ are unknown phases. In case of a perfect pulse, the transition probability becomes $p=1$ and $\epsilon=0$. However, this is often not the case, e.g., due to frequency or amplitude drifts or inhomogeneity, which make $\epsilon\neq 0$. For example, an amplitude error of $\Delta/\omega_{L}=(\omega_{L}-\Omega)/\omega_{L}=5\,\%$ from the target Rabi frequency of $\omega_{L}$ (Hartmann-Hahn resonance) leads to $\epsilon=1-\cos^{2}\left(\Delta t/2\right)\approx 0.06$ for a target $\omega_{L}t=\pi$ pulse, assuming $\delta=0$. Such errors can be compensated by applying phased sequences of pulses, where the phases of the subsequent pulses are chosen to cancel the errors of the individual pulses up to a certain order [35, 38, 45].
@@
and the phases $\phi_{k}$ of the individual pulses can be used as control parameters to achieve a robust performance. We can evaluate the latter by considering the fidelity [36, 38]